OpenStudy (anonymous):

Show that \[\sqrt{x+2}\rightarrow2 \] as \[x \rightarrow \infty\]

OpenStudy (anonymous):

I'll save you the trouble, it doesn't.

OpenStudy (australopithecus):

\[\lim_{x \rightarrow \infty} \sqrt{x+2} = \sqrt{\lim_{x \rightarrow \infty} x+2} = \sqrt{\infty + 2} = \sqrt{\infty} = \infty\] I can apply the rule \[\lim_{x \rightarrow \infty} f(x) = f(\lim_{x \rightarrow \infty}x)\] because the function is continuous look at the function 

OpenStudy (australopithecus):

it will keep increasing to infinity look at the graph

OpenStudy (anonymous):

Yeah sorry I got the question wrong.

OpenStudy (australopithecus):

no its ok, I hope my explanation was helpful